OpenStudy (anonymous):

solve the system of equations by substitution method x+7y=32 4x+3y=3

OpenStudy (anonymous):

y=5 x=-3

OpenStudy (anonymous):

how did u arrive at your answer mister anthonyanow?

OpenStudy (anonymous):

http://openstudy.com/code-of-conduct @anthonykanow @spimentel Solve the first equation for x and solve the second equation for y, then put the expression for x into the equation for y to get a value for y. Put that value of y into the equation for x and you'll have it.

OpenStudy (anonymous):

Ahh, I see. I apologize for my answer only response. Cliff is correct

OpenStudy (anonymous):

CAN YOU GUYS SHOW ME THE STEPS

OpenStudy (anonymous):

Start with the first equation x+7y=32 Solving for x, x=32-7y Now take that (32-7y) and put it in place of x in 4( )+3y=3